# 2]organization=ASML, city=Veldhoven, country=The Netherlands

### v1

@@ 3.1 Stage 1: Identification of additive MIMO systems
The identification problem is formulated based on the matrix residual, which is computed as the difference between the FRF measurement and the model, according to
@@ 3.2 Stage 2: Structured identification of modal models
Consider the nested model structures $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, where $\mathcal{M}_{1}$ denotes the modal model structure and $\mathcal{M}_{2}$ the additive model structure, with $\mathcal{M}_{1}\subset\mathcal{M}_{2}$. Since the model structures $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are nested, there exists a smooth and injective map $f:\mathcal{S}_{1}\to\mathcal{S}_{2}$, that links the parameter vectors $\boldsymbol{\rho}$ and $\boldsymbol{\beta}$ such that $\partial{f}({\boldsymbol{\rho}})/\partial{\boldsymbol{\rho}}^{\top}$ has full rank over $\mathcal{S}_{1}$. For the defined model structures, the subspace $\mathcal{S}_{1}=\mathbb{R}^{\operatorname{dim}(\boldsymbol{\rho})}$ consists of the vectors $\boldsymbol{\rho}$ as defined in (9), while the subspace $\mathcal{S}_{2}\subset\mathbb{R}^{\operatorname{dim}(\boldsymbol{\beta})}$ consists of the vectors $\boldsymbol{\beta}$ that admit a lower dimensional representation of the form $f(\boldsymbol{\rho})=\boldsymbol{\beta}$, i.e., the set of additive models satisfying the rank constraint. An estimate of the parameter vector $\boldsymbol{\rho}$ is then obtained as a solution to a weighted least-squares optimization problem
@@
Remark 2.
The two-stage approach for structured model estimation can be interpreted as a frequency-domain variant of the time-domain method introduced in Gonzalez2025StatisticallyIdentification , where it was shown that the method yields consistent and efficient estimates of structured model parameters in open loop, and consistent minimum-variance estimates in closed loop.
@@ 4.1 Initialization via linear least-squares
To solve the problem (21) an initial estimate of the model parameters $\boldsymbol{\beta}$ is needed. A convex least-squares method is introduced, which allows the numerator parameters to be computed in closed form, assuming fixed denominator polynomials. This reduces the initialization problem to determining initial pole locations, which are often effectively obtained from finite element models, FRF data, or tools such as complex mode indicator functions (CMIF) Shih1988ComplexEstimation . To this end, assume that the denominator polynomials are fixed at $\bar{A}_{i}(s)$, and let $\boldsymbol{\eta}$ represent the parameter vector from (18) without the denominator parameters. The estimate $\hat{\boldsymbol{\eta}}$ is found as the solution to the convex problem
@@
$\mathbf{\Phi}\left(\omega_{k},\boldsymbol{\beta}\right)=\left[\begin{array}[]{ccc}\mathbf{\Phi}_{1}^{\top}\left(\omega_{k},\boldsymbol{\theta}_{1}\right)&\ldots&\mathbf{\Phi}_{K}^{\top}\left(\omega_{k},\boldsymbol{\theta}_{K}\right)\end{array}\right]^{\top},\quad\mathbf{\Phi}_{i}\left(\omega_{k}\right)=\begin{bmatrix}\displaystyle\frac{\mathbf{I}_{n_{\mathrm{u}}n_{\mathrm{y}}}}{\xi^{\ell_{i}}\bar{A}_{i}(\xi_{k})}&\ldots&\displaystyle\frac{\xi^{m_{i}}\mathbf{I}_{n_{\mathrm{u}}n_{\mathrm{y}}}}{\xi^{\ell_{i}}\bar{A}_{i}(\xi_{k})}\end{bmatrix}^{\top}$. (24)
@@ 4.2.2 Derivation of the RIV estimator
Given an initial estimate $\boldsymbol{\beta}^{\langle 0\rangle}$ and a tolerance $\epsilon$, compute a new estimate until $\|\boldsymbol{\beta}^{\langle j+1\rangle}-\boldsymbol{\beta}^{\langle j\rangle}\|/\|\boldsymbol{\beta}^{\langle j\rangle}\|<\epsilon$ using:
@@ 5.1 Initial modal model via SVD decomposition
This decomposition is applied to the residue matrices estimated in the first stage, where the parameterization  (6) is considered for the general case, and  (11) for the proportionally damped case. The estimates of the real residue matrices $\hat{\mathbf{R}}_{i}$ are directly available, whereas the complex residue matrices $\hat{\mathbf{L}}_{i}$ are obtained by rearranging  (7) and (8), yielding
$\hat{\mathbf{L}}_{i}=\frac{1}{\boldsymbol{\lambda}_{i}-\bar{\boldsymbol{\lambda}}_{i}}\left(\hat{\mathbf{N}}_{i,0}+\boldsymbol{\lambda}_{i}\hat{\mathbf{N}}_{i,1}\right).$ (42)
The mode-shape vectors $\boldsymbol{\psi}_{l,i}$ and $\boldsymbol{\psi}_{r,i}$ are initialized using the dominant left and right singular vectors of the decomposition $\hat{\mathbf{L}}_{i}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\mathsf{H}}$. Similarly, the rigid-body mode-shape vectors $\boldsymbol{\phi}_{l,i}$ and $\boldsymbol{\phi}_{r,i}$ are obtained from the first $n_{\mathrm{rbm}}$ dominant singular vectors of the estimated matrix $\hat{\mathbf{R}}$, based on the SVD $\hat{\mathbf{R}}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\mathsf{H}}$. Note that the rigid-body modes require extraction of the first $n_{\mathrm{rbm}}$ singular vectors, as they are collectively represented by a single term in the additive model structure. For the proportional damping case, the mode-shape vectors are initialized following a similar procedure.
@@ 5.2 IPEM for modal model estimation
A key aspect of the IPEM problem (22) is the definition of the mapping $\mathbf{f}$, which relates the modal parameters in (9) to the additive parameters in (18). For the considered model structures, the mapping $\mathbf{f}$ is partitioned as
@@
where $\mathbf{g}$ maps the rigid-body mode parameters and each $\mathbf{h}_{i}$ maps the parameters of the $i$th flexible mode, from modal to additive representation. The specific form of these mappings depends on the assumed damping model and are provided in the subsequent subsections. A solution to the nonconvex problem in (22) is obtained using a Gauss–Newton (GN) method. At each iteration $j$, the residual
@@
Given an initial estimate $\boldsymbol{\rho}^{\langle 0\rangle}$ and tolerance $\epsilon$, compute a new estimate until $\|\boldsymbol{\rho}^{\langle j+1\rangle}-\boldsymbol{\rho}^{\langle j\rangle}\|/\|\boldsymbol{\rho}^{\langle j\rangle}\|<\epsilon$ by solving
@@
where $\alpha$ is a step-size parameter and $\mathbf{J}(\boldsymbol{\rho}^{\langle j\rangle})$ is the Jacobian matrix
@@
The converged solution of the iterations described by (45) yields a locally optimal solution to the original nonlinear least-squares problem (22). The following subsections provide the explicit forms of the parameter mappings, along with their associated Jacobians, for each of the considered model structures.
@@ 5.2.1 Parameter mappings for general-viscously damped mechanical systems
For the general damping model, the mapping functions $\mathbf{g}$ and $\mathbf{h}_{i}$ are derived considering the second-order parameterization (6), with the corresponding parameter vector defined in (9). The rigid-body modes are collectively represented by a single term in the additive model structure, hence, the function $\mathbf{g}$ defines a summation over each $n_{\mathrm{rbm}}$ residue matrices as follows
$\mathbf{g}(\boldsymbol{\varrho}_{1},\dots,\boldsymbol{\varrho}_{n_{\mathrm{rbm}}})=\operatorname{vec}\left(\sum_{i=1}^{n_{\mathrm{rbm}}}\boldsymbol{\phi}_{l,i}\boldsymbol{\phi}^{\top}_{r,i}\right).$ (47)
The function $\mathbf{h}_{i}$ for the $i$th flexible mode, with $i=1,\ldots,n_{\mathrm{flex}}$, is given as
$\mathbf{h}_{i}(\boldsymbol{\rho}_{i})=\begin{bmatrix}\boldsymbol{\lambda}_{i}\bar{\boldsymbol{\lambda}}_{i}&-\left(\boldsymbol{\lambda}_{i}+\bar{\boldsymbol{\lambda}}_{i}\right)&-\operatorname{vec}\left(\bar{\boldsymbol{\lambda}}_{i}\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}^{\top}_{r,i}+\boldsymbol{\lambda}_{i}\bar{\boldsymbol{\psi}}_{l,i}\bar{\boldsymbol{\psi}}^{\top}_{r,i}\right)^{\top}&\operatorname{vec}\left(\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}^{\top}_{r,i}+\bar{\boldsymbol{\psi}}_{l,i}\bar{\boldsymbol{\psi}}^{\top}_{r,i}\right)^{\top}\end{bmatrix}^{\top}$, (48)
where the first two entries relate the denominator coefficients and the last two entries the numerator matrices of each flexible mode. Since each mode is parameterized independently, the Jacobian in (46) takes a block-diagonal form
@@
$\mathbf{X}_{i}(\boldsymbol{\rho}_{i})=\begin{bmatrix}2\Re(\boldsymbol{\lambda}_{i})&2\Im(\boldsymbol{\lambda}_{i})&0&0&0&0\\ -2&0&0&0&0&0\\ 2\operatorname{vec}(\Re(\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}_{r,i}^{\top}))&-2\operatorname{vec}(\Im(\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}_{r,i}^{\top}))&\mathbf{X}_{i,33}&\mathbf{X}_{i,34}&\mathbf{X}_{i,35}&\mathbf{X}_{i,36}\\ 0&0&\mathbf{X}_{i,43}&\mathbf{X}_{i,44}&\mathbf{X}_{i,45}&\mathbf{X}_{i,46}\end{bmatrix}$, (51)
@@
$\begin{array}[]{ll}\begin{aligned} \mathbf{X}_{i,33}&=-\bar{\boldsymbol{\lambda}}_{i}(\boldsymbol{\psi}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}})-\boldsymbol{\lambda}_{i}(\bar{\boldsymbol{\psi}}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}}),\\ \mathbf{X}_{i,34}&=-j\bar{\boldsymbol{\lambda}}_{i}(\boldsymbol{\psi}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}})+j\boldsymbol{\lambda}_{i}(\bar{\boldsymbol{\psi}}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}}),\\ \mathbf{X}_{i,43}&=(\boldsymbol{\psi}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}})+(\bar{\boldsymbol{\psi}}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}}),\\ \mathbf{X}_{i,44}&=j(\boldsymbol{\psi}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}})-j(\bar{\boldsymbol{\psi}}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}}),\end{aligned}&\begin{aligned} \mathbf{X}_{i,35}&=-\bar{\boldsymbol{\lambda}}_{i}(\mathbf{I}_{n_{\mathrm{u}}}\otimes\boldsymbol{\psi}_{l,i})-\boldsymbol{\lambda}_{i}(\mathbf{I}_{n_{\mathrm{u}}}\otimes\bar{\boldsymbol{\psi}}_{l,i}),\\ \mathbf{X}_{i,36}&=-j\bar{\boldsymbol{\lambda}}_{i}(\mathbf{I}_{n_{\mathrm{u}}}\otimes\boldsymbol{\psi}_{l,i})+j\boldsymbol{\lambda}_{i}(\mathbf{I}_{n_{\mathrm{u}}}\otimes\bar{\boldsymbol{\psi}}_{l,i}),\\ \mathbf{X}_{i,45}&=(\mathbf{I}_{n_{\mathrm{u}}}\otimes\boldsymbol{\psi}_{l,i})+(\mathbf{I}_{n_{\mathrm{u}}}\otimes\bar{\boldsymbol{\psi}}_{l,i}),\\ \mathbf{X}_{i,46}&=j(\mathbf{I}_{n_{\mathrm{u}}}\otimes\boldsymbol{\psi}_{l,i})-j(\mathbf{I}_{n_{\mathrm{u}}}\otimes\bar{\boldsymbol{\psi}}_{l,i}).\end{aligned}\end{array}$
5.2.2 Parameter mappings for proportionally-damped mechanical systems
For proportionally damped systems, the mapping functions $\mathbf{g}$ and $\mathbf{h}_{i}$ are derived considering the second-order parameterization (6), with the corresponding parameter vector defined in (9). The rigid-body mode function $\mathbf{g}$ is identical as defined in (47), while the function $\mathbf{h}_{i}$ for the $i$th flexible mode, with $i=1,\ldots,n_{\mathrm{flex}}$, is given as
$\mathbf{h}_{i}\bigl(\boldsymbol{\rho}_{i}\bigr)=\begin{bmatrix}\omega_{i}^{2}&2\,\zeta_{i}\,\omega_{i}&\displaystyle\operatorname{vec}\bigl(\boldsymbol{\phi}_{l,i}\,\boldsymbol{\phi}_{r,i}^{\top}\bigr)^{\top}\end{bmatrix}$. (52)
Finally, the Jacobian matrix takes the same form as in (46), but with the Jacobian $\mathbf{X}_{i}$ of the $i$th flexible mode now computed according to

### v2

@@ 3.1 Stage 1: Identification of additive MIMO systems
The identification problem is formulated based on the matrix residual, which is computed as the difference between the FRF dataset $\{\mathbf{G}(\omega_{k})\}^{N}_{k=1}$ and the model, according to
@@ 3.2 Stage 2: Structured identification of modal models
Consider the nested model structures $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, where $\mathcal{M}_{1}$ denotes the modal model structure and $\mathcal{M}_{2}$ the additive model structure, with $\mathcal{M}_{1}\subset\mathcal{M}_{2}$. Since the model structures $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are nested, there exists a smooth and injective map $\mathbf{f}:\mathcal{S}_{1}\to\mathcal{S}_{2}$, that links the parameter vectors $\boldsymbol{\rho}$ and $\boldsymbol{\beta}$ such that $\partial{\mathbf{f}}({\boldsymbol{\rho}})/\partial{\boldsymbol{\rho}}^{\top}$ has full rank over $\mathcal{S}_{1}$. For the defined model structures, the subspace $\mathcal{S}_{1}=\mathbb{R}^{\operatorname{dim}(\boldsymbol{\rho})}$ consists of the vectors $\boldsymbol{\rho}$ as defined in (9), while the subspace $\mathcal{S}_{2}\subset\mathbb{R}^{\operatorname{dim}(\boldsymbol{\beta})}$ consists of the vectors $\boldsymbol{\beta}$ that admit a lower dimensional representation of the form $\mathbf{f}(\boldsymbol{\rho})=\boldsymbol{\beta}$, i.e., the set of additive models satisfying the rank constraint. The explicit form of $\mathbf{f}$ depends on the considered damping model and is further specified in Section 5. An estimate of the parameter vector $\boldsymbol{\rho}$ is then obtained as a solution to a weighted least-squares optimization problem
@@ 4.1 Initialization via linear least-squares
To solve the problem (21) an initial estimate of the model parameters $\boldsymbol{\beta}$ is needed. A convex linear least-squares method is introduced, which allows the numerator parameters to be computed in closed form, assuming fixed denominator polynomials. This reduces the initialization problem to determining initial pole locations, which are often effectively obtained from finite element models, FRF data, or tools such as complex mode indicator functions (CMIF) Shih1988ComplexEstimation . To this end, assume that the denominator polynomials are fixed at $\bar{A}_{i}(s)$, and let $\boldsymbol{\eta}$ represent the parameter vector from (18) without the denominator parameters. The estimate $\hat{\boldsymbol{\eta}}$ is found as the solution to the convex problem
@@
$\mathbf{\Phi}\left(\omega_{k}\right)=\left[\begin{array}[]{ccc}\mathbf{\Phi}_{1}^{\top}\left(\omega_{k}\right)&\ldots&\mathbf{\Phi}_{K}^{\top}\left(\omega_{k}\right)\end{array}\right]^{\top},\quad\mathbf{\Phi}_{i}\left(\omega_{k}\right)=\begin{bmatrix}\displaystyle\frac{\mathbf{I}_{n_{\mathrm{u}}n_{\mathrm{y}}}}{\xi^{\ell_{i}}\bar{A}_{i}(\xi_{k})}&\ldots&\displaystyle\frac{\xi^{m_{i}}\mathbf{I}_{n_{\mathrm{u}}n_{\mathrm{y}}}}{\xi^{\ell_{i}}\bar{A}_{i}(\xi_{k})}\end{bmatrix}^{\top}$. (24)
@@ 4.2.2 Derivation of the RIV estimator
Given an initial estimate $\boldsymbol{\beta}^{\langle 0\rangle}$ and tolerance $\epsilon_{1}$, compute a new estimate until $\|\boldsymbol{\beta}^{\langle j+1\rangle}-\boldsymbol{\beta}^{\langle j\rangle}\|/\|\boldsymbol{\beta}^{\langle j\rangle}\|<\epsilon_{1}$ using:
@@ 5.1 Initial modal model via SVD decomposition
This decomposition is applied to the residue matrices estimated in the first stage, where the parameterization  (6) is considered for the general case, and  (11) for the proportionally damped case. The estimates of the real residue matrices $\hat{\mathbf{R}}^{\mathrm{flex}}_{i}$ are directly available, whereas the complex residue matrices $\hat{\mathbf{L}}_{i}$ are obtained by rearranging  (7) and (8), yielding
$\hat{\mathbf{L}}_{i}=\frac{1}{\lambda_{i}-\bar{\lambda}_{i}}\left(\hat{\mathbf{N}}_{i,0}+\lambda_{i}\hat{\mathbf{N}}_{i,1}\right).$ (42)
The mode-shape vectors $\boldsymbol{\psi}_{l,i}$ and $\boldsymbol{\psi}_{r,i}$ are initialized using the dominant left and right singular vectors of the SVD decomposition $\hat{\mathbf{L}}_{i}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\mathsf{H}}$. In the additive model structure, all rigid-body modes are represented collectively by a single submodel, whereas in the modal representation they appear as a sum over $n_{\mathrm{rbm}}$ individual rigid-body modes. As a result, the corresponding residue matrix $\hat{\mathbf{R}}^{\mathrm{rbm}}\approx\sum_{j=1}^{n_{\mathrm{rbm}}}\mathbf{R}^{\mathrm{rbm}}_{j}$ contains the combined contribution of all rigid-body modes, with its dominant singular directions spanning the rigid-body subspace. Consequently, the initial estimates for the rigid-body mode-shape vectors $\boldsymbol{\phi}_{l,j}$ and $\boldsymbol{\phi}_{r,j}$ are obtained from the first $n_{\mathrm{rbm}}$ dominant singular vectors of the SVD $\hat{\mathbf{R}}^{\mathrm{rbm}}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\mathsf{H}}$. The mode-shape vectors for the proportional damping case are initialized following the same procedure.
@@ 5.2 IPEM for modal model estimation
A key aspect of the IPEM problem (22) is the definition of the function $\mathbf{f}:\boldsymbol{\rho}\mapsto\boldsymbol{\beta}$ which links the modal parameter vector (9) to the additive parameter vector (18). The function $\mathbf{f}$ specifies how a given set of modal parameters is translated into the corresponding coefficients of the additive model representation. In essence, $\boldsymbol{\rho}$ and $\mathbf{f}(\boldsymbol{\rho})=\boldsymbol{\beta}$ describe the same system but expressed in different parameterizations, and $\mathbf{f}$ provides the formal relationship between these two representations. For the considered model structures, the mapping $\mathbf{f}$ allows the partitioning
@@
where the function $\mathbf{g}:[\boldsymbol{\varrho}_{1}^{\top},\ldots,\boldsymbol{\varrho}_{n_{\mathrm{rbm}}}^{\top}]^{\top}\mapsto\boldsymbol{\theta}_{1}$ maps the rigid-body mode parameter vectors of the modal representation to the corresponding additive model parameter vector, and each $\mathbf{h}_{i}:\boldsymbol{\rho}_{i}\mapsto\boldsymbol{\theta}_{i}$ maps the parameter vector describing the $i$th flexible mode from modal representation to additive model representation. The explicit definition of the parameter mappings depend on the considered damping model and are provided in the subsequent subsections. A solution to the nonconvex problem in (22) can be obtained, e.g., using a Gauss–Newton (GN) procedure. At each iteration $j$, the residual
@@
Given an initial estimate $\boldsymbol{\rho}^{\langle 0\rangle}$ and tolerance $\epsilon_{2}$, compute a new estimate until $\|\boldsymbol{\rho}^{\langle j+1\rangle}-\boldsymbol{\rho}^{\langle j\rangle}\|/\|\boldsymbol{\rho}^{\langle j\rangle}\|<\epsilon_{2}$ by solving
@@
where $\alpha$ is a step-size parameter and $\mathbf{J}(\boldsymbol{\rho}^{\langle j\rangle})$ is the analytic Jacobian matrix evaluated at the current parameter estimate
@@
The converged solution of the iterations in (45) provides a locally optimal estimate of the nonlinear least-squares problem (22). Since the cost function is nonconvex, global optimality cannot be guaranteed and the obtained solution depends on the chosen initialization. The following subsections provide the explicit forms of the parameter mappings, along with their associated Jacobians, for each of the considered model structures.
@@ 5.2.1 Parameter mappings for general-viscously damped mechanical systems
For the general damping model, the mappings $\mathbf{g}$ and $\mathbf{h}_{i}$ are derived considering the second-order parameterization in (6), with the associated parameter vector defined in (9). The function $\mathbf{g}$ links the modal and additive parameterizations of the rigid-body modes, and since the additive representation describes all rigid-body dynamics using a single submodel, $\mathbf{g}$ amounts to summing the dyads of the $n_{\mathrm{rbm}}$ rigid-body mode-shape vectors according to
$\mathbf{g}(\boldsymbol{\varrho}_{1},\dots,\boldsymbol{\varrho}_{n_{\mathrm{rbm}}})=\operatorname{vec}\left(\sum_{j=1}^{n_{\mathrm{rbm}}}\boldsymbol{\phi}_{l,j}\boldsymbol{\phi}^{\top}_{r,j}\right).$ (47)
The function $\mathbf{h}_{i}$ links the modal and additive parameterizations of the $i$th flexible mode, with $i=1,\ldots,n_{\mathrm{flex}}$. The mapping includes the relation between eigenvalues and quadratic denominator parameters and the construction of the rank-2 numerator matrices from the complex mode-shape vectors. Specifically,
$\mathbf{h}_{i}(\boldsymbol{\rho}_{i})=\begin{bmatrix}\lambda_{i}\bar{\lambda}_{i}&-\left(\lambda_{i}+\bar{\lambda}_{i}\right)&-\operatorname{vec}\left(\bar{\lambda}_{i}\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}^{\top}_{r,i}+\lambda_{i}\bar{\boldsymbol{\psi}}_{l,i}\bar{\boldsymbol{\psi}}^{\top}_{r,i}\right)^{\top}&\operatorname{vec}\left(\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}^{\top}_{r,i}+\bar{\boldsymbol{\psi}}_{l,i}\bar{\boldsymbol{\psi}}^{\top}_{r,i}\right)^{\top}\end{bmatrix}^{\top}$, (48)
where the first two entries correspond to the denominator coefficients, and the last two entries correspond to the numerator matrices. Since each mode is parameterized independently, the Jacobian in (46) takes a block-diagonal form
@@
$\mathbf{X}_{i}(\boldsymbol{\rho}_{i})=\begin{bmatrix}2\Re(\lambda_{i})&2\Im(\lambda_{i})&0&0&0&0\\ -2&0&0&0&0&0\\ 2\operatorname{vec}(\Re(\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}_{r,i}^{\top}))&-2\operatorname{vec}(\Im(\boldsymbol{\psi}_{l,i}\boldsymbol{\psi}_{r,i}^{\top}))&\mathbf{X}_{i,33}&\mathbf{X}_{i,34}&\mathbf{X}_{i,35}&\mathbf{X}_{i,36}\\ 0&0&\mathbf{X}_{i,43}&\mathbf{X}_{i,44}&\mathbf{X}_{i,45}&\mathbf{X}_{i,46}\end{bmatrix}$, (51)
@@
$\begin{array}[]{ll}\begin{aligned} \mathbf{X}_{i,33}&=-\bar{\lambda}_{i}(\boldsymbol{\psi}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}})-\lambda_{i}(\bar{\boldsymbol{\psi}}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}}),\\ \mathbf{X}_{i,34}&=-j\bar{\lambda}_{i}(\boldsymbol{\psi}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}})+j\lambda_{i}(\bar{\boldsymbol{\psi}}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}}),\\ \mathbf{X}_{i,43}&=(\boldsymbol{\psi}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}})+(\bar{\boldsymbol{\psi}}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}}),\\ \mathbf{X}_{i,44}&=j(\boldsymbol{\psi}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}})-j(\bar{\boldsymbol{\psi}}_{r,i}\otimes\mathbf{I}_{n_{\mathrm{y}}}),\end{aligned}&\begin{aligned} \mathbf{X}_{i,35}&=-\bar{\lambda}_{i}(\mathbf{I}_{n_{\mathrm{u}}}\otimes\boldsymbol{\psi}_{l,i})-\lambda_{i}(\mathbf{I}_{n_{\mathrm{u}}}\otimes\bar{\boldsymbol{\psi}}_{l,i}),\\ \mathbf{X}_{i,36}&=-j\bar{\lambda}_{i}(\mathbf{I}_{n_{\mathrm{u}}}\otimes\boldsymbol{\psi}_{l,i})+j\lambda_{i}(\mathbf{I}_{n_{\mathrm{u}}}\otimes\bar{\boldsymbol{\psi}}_{l,i}),\\ \mathbf{X}_{i,45}&=(\mathbf{I}_{n_{\mathrm{u}}}\otimes\boldsymbol{\psi}_{l,i})+(\mathbf{I}_{n_{\mathrm{u}}}\otimes\bar{\boldsymbol{\psi}}_{l,i}),\\ \mathbf{X}_{i,46}&=j(\mathbf{I}_{n_{\mathrm{u}}}\otimes\boldsymbol{\psi}_{l,i})-j(\mathbf{I}_{n_{\mathrm{u}}}\otimes\bar{\boldsymbol{\psi}}_{l,i}).\end{aligned}\end{array}$
5.2.2 Parameter mappings for proportionally damped mechanical systems
For proportionally damped systems, the mapping functions $\mathbf{g}$ and $\mathbf{h}_{i}$ are derived considering the second-order parameterization in (6), with the associated parameter vector defined in (9). The rigid-body mapping $\mathbf{g}$ is identical to the expression in (47). For the $i$th flexible mode, $i=1,\ldots,n_{\mathrm{flex}}$, the function $\mathbf{h}_{i}$ takes the form
$\mathbf{h}_{i}\bigl(\boldsymbol{\rho}_{i}\bigr)=\begin{bmatrix}\omega_{i}^{2}&2\,\zeta_{i}\,\omega_{i}&\displaystyle\operatorname{vec}\bigl(\boldsymbol{\phi}_{l,i}\,\boldsymbol{\phi}_{r,i}^{\top}\bigr)^{\top}\end{bmatrix}$, (52)
where the first two entries define the denominator coefficients, while the final entry contains the numerator terms constructed from the real mode-shape vectors. Finally, the Jacobian matrix takes the same form as in (46), but with the Jacobian $\mathbf{X}_{i}$ of the $i$th flexible mode now computed according to